OpenStudy (anonymous):

Determine whether each equation has the given ordered pair as a solution. 5x − 3y = 15; (0, 5)

OpenStudy (oaktree):

Plug the numbers in. Does it work?

OpenStudy (oaktree):

5(5)-3(0) = 25-0 = 25. Does it?

OpenStudy (oaktree):

No, because 25 is not 15. So, no.

OpenStudy (anonymous):

ok